OpenStudy (anonymous):

In the following system of equations, what does x equal? -2x – 0y = -4 -6x – 2y = -12

OpenStudy (anonymous):

does it really say \[-2x-0y=-4\]???

OpenStudy (anonymous):

who writes this stuff? \[-2x=-4\] \[x=2\] \[-6\times 2-2y=-12\] \[-12-2y=-12\] \[-2y=0\] \[y=0\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its stupid summer school stuff

OpenStudy (anonymous):

stupid doesn't begin to describe it. is it some on line machine that writes \[-2x-0y=-4\]? no human being would do it